SpaceRef

SpaceRef


SOHO Daily Report Operations Day - 070/2003 (11 March 03)

Status Report From: Goddard Space Flight Center
Posted: Thursday, March 13, 2003

Mission Day 2657

SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC

Ref. #: SH-FOT-DR-030311

PASS #1 Madrid DSS54 0915-1841

PASS #2 Goldstone DSS27 2137-071/0205

PASS #1 Madrid DSS54 0915-1841

Anomalies:

G03-0171 0930 Degraded TM; Cause unknown (550 mins, recov.) DR #N400851

G03-0172 1608 Two wayward commands during CDS NRT (17 min NRT delay) DR #N400852, DR #N400853

NRT:

1529-1747

Activities Summary:

  • - SSR Dump
  • - Collected MDI-M (0942-0958, 1240-1820)
  • - Uplinked delayed loads (VIRA30312, SWA030311101)
  • - Reaction Wheel Speed Monitoring Daily Adjustment
  • - CRS Drift Estimation Update
  • - Uplinked RSL Table
  • - Executed CEPAC Contingency Script H-09 (ESU Reset) per OCD# 1764
  • - New Guide Star (5.0 mag, slot 3)
  • - SSR SEF Reset (x2)
  • - SSR Record

PASS #2 Goldstone DSS27 2137-071/0205

Anomalies:

G03-0173 2135 Late AOS due to antenna azimuth drive problem (2 mins, covered by SSR) DR #G102464

G03-0174 2142 Bad timestamps on TM packets; TCA problem (6 mins) DR #G102465

NRT:

None

Activities Summary:

  • - SSR Dump
  • - Collected MDI-M (2310-0145)
  • - Uplinked delayed load (CEL030311301)
  • - OBT Distribution to SWAN, CELIAS
  • - SSR SEF Reset
  • - SSR Record

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks